OpenStudy (anonymous):

Evaluate the function at the given value of the independent variable and simplify. f(x) = 3x2 - 4x + 4; f(x - 1) A) 3x2 - 10x + 11 B) -10x2 + 3x + 11 C) 3x2 + 8x + 3 D) 3x2 - 10x + 3

OpenStudy (anonymous):

start with \[f(x-1)=3(x-1)^2-4(x-1)+4\] and then do a raft of algebra

OpenStudy (anonymous):

3x\[3x ^{2}-3^{2}-4x-4+4\]

OpenStudy (anonymous):

@satellite73 is that correct or is its +4+4

OpenStudy (anonymous):

i think there isa mistake there

OpenStudy (anonymous):

when you multiply out, the first term is \[3(x-1)^2=3(x^2-2x+1)=3x^2-6x+3\]

OpenStudy (anonymous):

second term is \(- 4(x-1)=-4x+4\)

OpenStudy (anonymous):

so you should get \[3x^2-6x+3-4x+4+4\] then combine like terms remember i said "do a raft of algebra" don't make up any shortcuts

OpenStudy (anonymous):

so is it 3x^2-10x+11

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

ty i have tons more problems if you like to help